# MNSC-0001-2024.00

### v1

@@ 4.1 System Model
Fig. 4 depicts the four key roles within AIGC-Chain: Producer, Provider, Consumer, and Auditor. These roles interact with AIGC-Chain by submitting transactions to record copyright information on the blockchain, facilitating secure management of AIGC copyright details. Furthermore, the system incorporates a Fabric CA, which is tasked with providing digital certification services to users participating in the system, streamlining their involvement in blockchain-based activities. Here are the detailed descriptions of the four roles:
@@
  • Producer: Producer is the initiator of AIGC generation, existing as a light node on AIGC-Chain. It selects an appropriate model and inputs a series of parameters and keywords to trigger the generation. Producer is the initial owner of the AIGC product, with the right to use, share, and transfer its AIGC product.
  • Provider: Provider owns the trained AIGC model and exists as a full node on AIGC-Chain. It can provide AIGC generation service to Producer and charge a fee. Provider is responsible for promptly recording the usage and generation activities of models on AIGC-Chain.
@@ 4.2 Threat Model
The threats considered in our scheme are as follows, which come from four entities, i.e., Producer, Provider, Consumer.
  • Producer, Provider, and Consumer may intentionally submit false or misleading proof to claim copyright ownership over the AIGC product.
  • Producer and Provider might furnish the Auditor with erroneous or fabricated transaction ID, which could lead to incorrect copyright assessments for AIGC products, thereby misleading the Auditor and preventing the tracing of the accurate copyright information related to the AIGC products.
@@ 4.3 Design Goals
  • Reliability: Throughout each phase of AIGC creation, the intermediate data outputs are meticulously and accurately uploaded to the AIGC chain. By harnessing the inherent immutability of blockchain, we guarantee the accuracy and reliability of the copyright proof retrieved, facilitating the management of copyright ownership.
  • Traceability: All data generated throughout the full lifecycle of an AIGC product can be traced and verified comprehensively and accurately, with no instances of fabrication or alteration of copyright-related data.
@@ 5 The Proposed System
In this section, we delve into the specifics of AIGC-Chain. We begin by providing a comprehensive system overview. Subsequently, we introduce a novel approach for tracing data within the blockchain. We conclude by elaborating on the operational workflow and the corresponding state transitions within the proposed system. For improved readability, a concise compilation of the key notations used in this section, including their definitions, is provided in Table 1.
Table 1: NOTATIONS AND DESCRIPTIONS
@@
Producer Sign the signature of Producer
@@ 5.1 Overview
By storing data generated throughout the AIGC lifecycle in a complete and unchangeable form on a distributed blockchain, we can ensure the traceability of the sources of AIGC products. This also guarantees that in the event of copyright disputes, data stored on AIGC-Chain can be used to determine the ownership of AIGC products, thereby protecting the security of AIGC copyrights. As depicted in Fig. 4, AIGC-Chain comprises 5 core procedures: Registration, Content Generation, Data Uploading, Copyright Trading, and Copyright Management.
AIGC-Chain primarily relies on a consortium chain implemented on Hyperledger Fabric. It maintains its data in a decentralized blockchain ledger composed of blocks and the world state. Each block contains records of all preceding transactions that have been successfully executed, while the world state preserves the latest data state in the form of key-value pairs.
Within AIGC-Chain, a world state of 8 items is preserved in the ledger of AIGC-Chain for each AIGC product, as illustrated in Fig. 4. This world state undergoes continuous updates throughout the lifecycle of the AIGC, responsive to a multitude of operations. Among them, Model ID serves as a unique identifier for the training models employed. Given that participants on a consortium chain have full access to all data on the chain, we employ a strategy to protect the privacy of the AIGC product. Instead of storing the original content, we use a hash value Description Hash, derived from the input Args to update the world state. Model Sign and Producer Sign respectively denote the signatures of the model owner and the producer of the AIGC product. Once established, these signatures are permanent and cannot be modified. Similarly, Owner Sign is the signature of the owner of the AIGC product. Before the AIGC product is sold, the owner and the producer are identical. Following the transaction, this attribute will be updated. Status records the current state of the AIGC product, such as generating, generated, etc. Additionally, BF is employed to index all transaction information related to the AIGC product on the blockchain, facilitating the authenticity management of copyright. Details of BF will be discussed in Section 5.2.
@@ 5.2 Efficient Tracing Method Based on Indistinguishable Bloom Filter
Each product within the AIGC-Chain is associated with numerous transactions spread across various blocks. To acquire the most up-to-date information, one can retrieve it from the world state of the ledger. Meanwhile, for historical data, examining the transaction specifics within these blocks is necessary. In copyright disputes on the AIGC-Chain, auditors can refer to the world state using Product ID to access product details and can track more granular historical generation information through Txids provided by various participants, including timestamps and content alterations. However, since blockchain is publicly accessible to all participants, this may potentially allow malicious actors to interfere with the judgments of auditors by submitting similar transactions to impersonate copyright owners.
Fast Query (Chen et al. 2023c) was proposed to enhance the retrieval speed of copyright-related transactions on the blockchain. It stores a table that contains all the copyright in the blockchain according to the hash value of the copyright and a one-way linked list for each copyright in the table, allowing for queries to be performed solely on the locally stored data, obviating the need to traverse all transaction information stored on the chain. While this design significantly improves the efficiency of information retrieval on the blockchain, it also introduces additional offline storage expenses and security concerns.
To counter this, we propose an efficient copyright tracing method IBFT based on the Indistinguishable Bloom Filter (IBF), which adds an array BF of $m$ twins for each product in the world state to map all related transactions. IBF can obfuscate values within the Bloom Filter, thereby preserving the privacy of the data. This approach not only effectively deters fraudulent behavior in copyright disputes but also streamlines the audit process for AIGC copyrights.
The data structure of the transaction in Fabric is depicted in Fig. 5. Within this structure, Args contains the input parameter for the transaction, while CreateTime denotes the timestamp of transaction creation. Although malicious attackers might forge transactions by invoking transactions with the same Args, they are unable to fabricate a new transaction that is identical in critical parameters such as CreateTime and Nonce to the original. Consequently, we propose mapping Args along with CreateTime into BF in the world state. Only transactions with a valid Txid can pass the verification of IBFT, thereby effectively reducing the risk of copyright tampering.
IBFT is composed of smart contracts that are automatically executed upon the triggering of transactions. As presented in Algorithm 1, it consists of three procedures, which are described in detail as follows:
@@
1) Setup: Upon receiving an AIGC generation request from Producer, an array BF and hash functions $\left\{H_{key},H\right\}$ are generated specifically for product P. They are then stored within the transaction and world state on the chain, with access restricted to Provider, Consumers, and Auditors associated with product P.
@@
2) Insert: Upon initiating a transaction related to product P, BF in the world state will be updated as:
$BF\left[(h_{i}(v))\right]\left[H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=1$
$BF\left[(h_{i}(v))\right]\left[1-H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=0$
@@
where $v=Args||CreateTime$ is retrieved from the proposed transaction Tx, and $i\in[1,k]$, $k$ represents the number of hash mapping functions..
@@
3) Check: To confirm the authenticity of the transaction $Tx^{\prime}$ received from participants, check $v^{\prime}=Args^{\prime}||CreateTime^{\prime}$ within $Tx^{\prime}$ as:
$BF\left[(h_{i}(v^{{}^{\prime}}))\right]\left[H(h_{k+1}(h_{i}(v^{{}^{\prime}}))\oplus\gamma\right]=1$
@@
If $v^{\prime}$ of $k$ selected cells for BF is found to be 1, it indicates that $Tx^{\prime}$ is linked to the generation of the AIGC product P.
Algorithm 1 Details of IBFT
@@
$v=Args^{\prime}||CreateTime^{\prime}$;
@@ 5.3 Lifecycle Management of AIGC Product
We incorporate Producer Sign into the world state of AIGC products, establishing a direct connection between each product and the personal information of producers. This enhances the traceability of the AIGC products, allowing for the efficient monitoring of user identity information, including that of the producer. As a result, the measure effectively prevents the misuse of AIGC models and significantly reduces the production of fraudulent or low-quality AIGC content. To prevent duplicate transactions of product copyrights, we add Owner Sign in the world state of AIGC products, which ensures that each copyright transaction is permanently recorded on the blockchain. Consumers can review the transaction history of the product and modify Owner Sign upon completion of the deal. The transaction process is contingent upon the validation of the ownership signature; only when the signature is successfully verified can the transaction proceed.
@@
In Registration, CA registers roles participating in AIGC-Chain as lightweight or full nodes. Content Generation takes the prompt provided by Producer to the system and assigns unique Product ID as key for the aigproduct P. Instead of storing the actual prompt, it saves the hash of the prompt as Description Hash in the world state. After completing the content generation, Provider sends the hash value and other related generation information of P through Data Uploading and returns P to Producer. Consumer and Producer execute copyright transactions and ownership transfers for AIGC product through Copyright Trading. In the event of a copyright dispute, Auditors can invoke Copyright Management to retrieve the information related to P from AIGC-Chain. Following this, the auditors can proceed to analyze and evaluate the copyright of the product. The detailed introductions of the above five stages are as follows:
@@
Within this architecture, Producer and Consumer participate as lightweight nodes, predominantly hosted on personal computers and mobile devices. They can engage with blockchain interactions without being responsible for data storage. In contrast, Provider and Auditor, armed with more robust computing and storage power, serve as full nodes within the network, tasked with the responsibility of synchronizing the entire data on the blockchain.
@@
Producer:
$\sigma_{1}=sig_{sk_{Producer}}(Args)$;
@@
Verify $\sigma_{1}$ with $pk_{Producer}$;
@@
Send $\sigma_{2}$ back to Producer;
Producer:
@@
Owner $\textit{Owner Sign}=sig_{sk_{Producer}}(Description~{}Hash)$;
@@
2) Content Generation: When using an AIGC model to generate product P, Producer inputs the prompt $Args$ into the model and selects a unique Product ID as key for P. Next, Producer calls the smart contract to initiate a transaction $Tx_{req}$ to update the world state for P with key. As illustrated in Fig. 6, transaction $Tx_{req}$ updates the world state by updating the Description Hash, Producer Sign, Owner Sign, BF, and Status. Since the world state is public to all participants on the consortium chain, we use the hash of the prompt Description Hash in $Tx_{req}$ to prevent the prompt from being copied or maliciously used by an attacker. The specific process of this stage is as follows:
  • Producer computes $\sigma_{1}=sig_{sk_{Producer}}(Args)$ and sends $\sigma_{1}$ along with $Args$ to Provider for content generation.
  • Provider verifies $\sigma_{1}$ with $pk_{Producer}$ of Producer. The successful verification suggests that the prompt $Args$ aligns with the intent of the Producer and has not been altered during transmission.
  • Provider generates $\sigma_{2}=sig_{sk_{Provider}}(Args)$ and sends it back to Producer. This indicates that Provider acknowledges the legitimacy of the received $Args$, enabling further content generation operations.
  • Producer selects a random hash function $H$ and computes Description $\textit{Description ID}=H(Args)$.
  • Producer verifies the legality of $\sigma_{2}$; a successful validation implies the formation of consensus between Producer and Provider.
  • Producer generates signature Owner Sign with Description Hash.
  • A transaction $Tx_{req}$ is proposed by Producer to update the world state with $key=\textit{Product ID}$, where Status is used to indicate the current state of P, which is prepared.
  • Producer retrieves the transaction ID $Txid_{req}$ from transaction $Tx_{req}$ and may choose to store it either locally or within a decentralized storage system such as IPFS [^5] https://www.ipfs.tech, which facilitates the simplification of the copyright management process.
@@
Send $\sigma_{3}$ and generated product P back to Producer;
Producer:
@@
Verify the legality of $\sigma_{4}$ with $pk_{Producer}$;
@@
3) Data Uploading: Provider generates product P with the AIGC model based on the prompt $Args$ provided by Producer. Subsequently, Provider transmits P to Producer and initiates a transaction $Tx_{gen}$ to update additional information related to the generation process in the world state. Upon receiving product P, Producer initiates transaction $Tx_{upload}$ to add product P into the world state. Fig. 6 illustrates the alteration in the world state following the execution of transactions $Tx_{gen}$ and $Tx_{upload}$. In this process, the use of Product Hash in place of the original product P is aimed at preventing potential copyright infringement or misuse, while also ensuring the security of the data in AIGC-Chain.
@@
  • Provider sends both $\sigma_{3}$ and the generated product P back to Producer for verification.
  • Producer authenticates the validity of $\sigma_{3}$; only upon successful verification will Producer accept product P and proceed to sign it as $\sigma_{4}=sig_{sk_{Producer}}(P)$, which will be sent back to Provider.
  • Provider verifies the legality of $\sigma_{4}$; a successful validation implies the formation of consensus between Producer and Provider.
@@
  • Producer computes Product $\textit{Product Hash}=H(P)$ for product P.
  • A transaction $Tx_{upload}$ is proposed by Producer to update the world state, where Status has transitioned from generating to generated.
  • Producer and Provider retrieve transaction ID $Txid_{gen}$ and $Txid_{upload}$ from the blockchain respectively and can choose to store them either locally or within a decentralized storage system such as IPFS.
@@
Figure 7: The Process of Copyright Management
5) Copyright Management: For a copyright dispute, Auditor can retrieve Product ID of the disputed product P by consulting the relevant world state on the AIGC-Chain. As depicted in Fig. 7, Provider and Provider are required to provide proof to Auditor to establish their connection with the disputed product. Auditor will then assess the ownership of the copyright based on the proof gathered.
  • Producer submits the proof $\varphi$ to Auditor, where $\varphi=(pk_{Producer},Txid_{req},Txid_{upload},prompt)$.
@@
  • Auditor utilizes $pk_{Producer}$ and $pk_{Provider}$ to verify Producer Sign and Model Sign retrieved from the world state of the AIGC-Chain for Product ID, to assess the correctness of the identities associated with the product-related roles.
  • Auditor employs IBFT.Check to verify the correctness of $Txid_{req}$, $Txid_{upload}$, and $Txid_{gen}$. An output of 1 indicates that the transaction is indeed related to the generation process of the AIGC product.
  • Auditor conducts an in-depth manual review of the originality, compliance, ownership, and value of the AIGC product based on the information provided by Producer and Provider, as well as the product information stored on AIGC-Chain.

### v2

@@ 4.1 System Model
Fig. 4 depicts the four key roles within AIGC-Chain: User, Provider, Consumer, and Auditor. These roles interact with AIGC-Chain by submitting transactions to record copyright information on the blockchain, facilitating secure management of AIGC copyright details. Here are the detailed descriptions of the four roles:
@@
  • User: User is the initiator of AIGC generation, existing as a light node on AIGC-Chain. It selects an appropriate model and inputs a series of parameters and keywords to trigger the generation. User is the initial owner of the AIGC product, with the right to use, share, and transfer its AIGC product.
  • Provider: Provider owns the trained AIGC model and exists as a full node on AIGC-Chain. It can provide AIGC generation service to User and charge a fee. Provider is responsible for promptly recording the usage and generation activities of models on AIGC-Chain.
@@ 4.2 Threat Model
The threats considered in our scheme are as follows, which come from four entities, i.e., User, Provider, Consumer.
  • User, Provider, and Consumer may intentionally submit false or misleading proof to claim copyright ownership over the AIGC product.
  • User and Provider might furnish the Auditor with erroneous or fabricated transaction ID, which could lead to incorrect copyright assessments for AIGC products, thereby misleading the Auditor and preventing the tracing of the accurate copyright information related to the AIGC products.
@@ 4.3 Design Goals
  • Trustworthy: Throughout each phase of generation, the intermediate data outputs are meticulously and accurately uploaded to the AIGC-Chain. By harnessing the inherent immutability of blockchain, we guarantee the trustworthy of the copyright proof retrieved, facilitating the management of copyright ownership.
  • Traceability: All data generated throughout the full lifecycle of an AIGC product can be permanently recorded and cryptographically verifiable, eliminating risks of tampering or unauthorized modifications to copyright-related information.
@@ 5 The Proposed System
In this section, we delve into the specifics of the proposed system. We begin by providing a comprehensive overview. Subsequently, we introduce a novel approach for tracing data within the blockchain. We conclude by elaborating on the operational workflow and the corresponding state transitions within the proposed system. For improved readability, a concise compilation of the key notations used in this section, including their definitions, is provided in Table I.
TABLE I: NOTATIONS AND DESCRIPTIONS
@@
User Sign the signature of User
@@ 5.1 Overview
By storing data generated throughout the AIGC lifecycle in a complete and unchangeable form on a distributed blockchain, we can ensure the trustworthy and traceability of the sources of AIGC products. This also guarantees that in the event of copyright disputes, data stored on AIGC-Chain can be used to determine the ownership of AIGC products, thereby protecting the security of AIGC copyrights. As depicted in Fig. 4, AIGC-Chain comprises 5 core procedures: Registration, Content Generation, Data Uploading, Copyright Trading, and Copyright Management.
AIGC-Chain is deployed on a private Ethereum network, which maintains its data in a decentralized blockchain ledger composed of blocks and the world state. Each block contains records of all preceding transactions that have been successfully executed, while the world state preserves the latest data state in the form of key-value pairs.
Within AIGC-Chain, a world state of 8 items is preserved in the ledger of AIGC-Chain for each AIGC product, as illustrated in Fig. 4. This world state undergoes continuous updates throughout the lifecycle of AIGC products, responsive to a multitude of operations. Among them, Model ID serves as a unique identifier for the training models employed. Given that participants of the blockchain have full access to all data on the chain, we employ a strategy to protect the privacy of AIGC products. Instead of storing the original content, we use a hash value Description Hash, derived from the input Args to update the world state. Model Sign and User Sign respectively denote the signatures of the model owner and the producer of the AIGC product. Once established, these signatures are permanent and cannot be modified. Similarly, Owner Sign is the signature of the owner of the AIGC product. Before the AIGC product is sold, the owner and the producer are identical. Following the transaction, this attribute will be updated. Status records the current state of the AIGC product, such as generating, generated, etc. Additionally, BF is employed to index all transaction information related to the AIGC product on the blockchain, facilitating the authenticity management of copyright. Details of BF will be discussed in Section 5.2.
@@ 5.2 Efficient Tracing Method Based on Indistinguishable Bloom Filter
As illustrated in Fig. 6, the generation of each AIGC product is inherently associated with at least three categories of on-chain transactions. In the event of a copyright dispute, Auditors could access detailed copyright proof through Product ID and require all participants to submit copyright proofs of their involvement in the generation process, enabling fine-grained copyright traceability. It should be noted that while the input parameters Args in transactions can validate the authenticity of transactions, the transparent nature of blockchain allows potential attackers to forge transactions with identical Args in attempts to manipulate copyright attribution of AIGC products. Chen et al. [9] proposed a fast query algorithm based on a singly-linked list, which constructs a local hash index structure to store the hash values of copyright information in a singly linked list. Compared with the traditional linear retrieval approach requiring full-node traversal, this algorithm enables copyright ownership verification without parsing the entire blockchain transaction data. However, this centralized storage system necessitates continuous maintenance of offline index replicas, leading to increased local storage costs. More critically, it also brings the risk of a single point of failure.
To address this vulnerability, we propose CTtrace, an efficient copyright traceable verification mechanism based on indistinguishable bloom filters. It constructs an $m$-bit twin array BF for each blockchain-recorded AIGC product with all associated transactions mapped. By implementing hash mapping of input parameter Args, timestamp CreateTime, and random value Nonce—as depicted in Fig. 5CTtrace ensures that even if attackers duplicate Args, they cannot reconstruct legitimate transaction records with identical timestamps and nonces.
Specifically, as depicted in Algorithm 1, CTrace comprises three stages: Setup, Insert, and Check. The smart contract automatically executes these stages upon triggering transactions for copyright storage and querying.
@@
1) Setup: Upon the receipt of an AIGC generation request from User, a dedicated array BF and a set of hash functions $\left\{H_{key},H\right\}$ are generated for product P. These elements are subsequently embedded within the transaction and recorded in the world state on the blockchain. Access to this information is granted exclusively to Provider, Consumers, and Auditors affiliated with product P.
@@
2) Insert: When a transaction is initiated to update the status of product P, its BF in the world state will be revised to:
$BF\left[(h_{i}(v))\right]\left[H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=1,$
$BF\left[(h_{i}(v))\right]\left[1-H(h_{k+1}(h_{i}(v))\oplus\gamma\right]=0,$
@@
where $v=Args||CreateTime||Nounce$ is retrieved from the proposed transaction Tx, and $i\in[1,k]$, $k$ represents the number of hash mapping functions.
@@
3) Check: To confirm the authenticity of the transaction $Tx^{\prime}$ received from participants, check $v^{\prime}=Args^{\prime}||CreateTime^{\prime}||Nounce^{\prime}$ within $Tx^{\prime}$ as:
$BF\left[(h_{i}(v^{{}^{\prime}}))\right]\left[H(h_{k+1}(h_{i}(v^{{}^{\prime}}))\oplus\gamma\right]=1.$
@@
If $v^{\prime}$ of $k$ selected hash functions for BF is found to be 1, it indicates that $Tx^{\prime}$ is linked to the generation of product P.
Algorithm 1 Details of CTrace
@@
  $v=Args^{\prime}||CreateTime^{\prime}||Nounce^{\prime}$;
@@
For instance, in the content generation based on diffusion models, the core workflow can be decomposed into two sequential phases: Forward Diffusion and Reverse Denoising, which collaboratively accomplish content generation through a multi-round iterative process. During the Forward Diffusion phase, key parameters are immortalized on the blockchain, encompassing noise schedule ${\beta_{1},\beta_{2},\ldots,\beta_{T}}$, the set of noisy images ${x_{1},x_{2},\ldots,x_{T}}$, and noise samples $\epsilon$ that are introduced. In the subsequent Reverse Denoising phase, blockchain records the set of denoised images ${x_{T},x_{T-1},\ldots,x_{0}}$ and the set of evolving parameter for the predicted noise $\epsilon_{\theta}(x_{t},t)$. Furthermore, considering the request transaction $Tx_{req}$ and multiple content exchange transactions $Tx_{exchange}$, we assume that the number of associated transactions per generated product is constrained by $n\leq 20$. To address the correctness verification requirements, when the false positive rate threshold is set to 0.1%, probability model derivation yields the optimal parameter combination $(k=10,m=288)$ satisfying:
$P_{\textit{fp}}\approx\left(1-e^{-kn/m}\right)^{k}\leq 0.001,$
This configuration guarantees that the system upholds the necessary probabilistic assurances, striking a balance between computational efficiency and the reliability of verification.
@@ 5.3 Lifecycle Management of AIGC Product
We incorporate User Sign into the world state of AIGC products, establishing a direct connection between each product and the personal information of producers. This enhances the traceability of the AIGC products, allowing for the efficient monitoring of user identity information, including that of the producer. As a result, the measure effectively prevents the misuse of AIGC models and significantly reduces the production of fraudulent or low-quality AIGC content. To prevent duplicate transactions of product copyrights, we add Owner Sign in the world state of AIGC products, which ensures that each copyright transaction is permanently recorded on the blockchain. Consumers can review the transaction history of the product and modify Owner Sign upon completion of the deal. The transaction process is contingent upon the validation of the ownership signature; only when the signature is successfully verified can the transaction proceed.
@@
In Registration, CA registers roles participating in AIGC-Chain as lightweight or full nodes. Content Generation takes the prompt provided by User to the system and assigns unique Product ID as key for the aigproduct P. Instead of storing the actual prompt, it saves the hash of the prompt as Description Hash in the world state. After completing the content generation, Provider sends the hash value and other related generation information of P through Data Uploading and returns P to User. Consumer and User execute copyright transactions and ownership transfers for AIGC product through Copyright Trading. In the event of a copyright dispute, Auditors can invoke Copyright Management to retrieve the information related to P from AIGC-Chain. Following this, the auditors can proceed to analyze and evaluate the copyright of the product. The detailed introductions of the above five stages are as follows:
@@
Within this architecture, User and Consumer participate as lightweight nodes, predominantly hosted on personal computers and mobile devices. They can engage with blockchain interactions without being responsible for data storage. In contrast, Provider and Auditor, armed with more robust computing and storage power, serve as full nodes within the network, tasked with the responsibility of synchronizing the entire data on the blockchain.
@@
  User:
  $\sigma_{1}=sig_{sk_{User}}(Args)$;
@@
  Verify $\sigma_{1}$ with $pk_{User}$;
@@
  Send $\sigma_{2}$ back to User;
  User:
@@
  Owner $\textit{Owner Sign}=sig_{sk_{User}}(Description~{}Hash)$;
@@
2) Content Generation: When using an AIGC model to generate product P, User inputs the prompt $Args$ into the model and selects a unique Product ID as key for P. Next, User calls the smart contract to initiate a transaction $Tx_{req}$ to update the world state for P with key. As illustrated in Fig. 6, transaction $Tx_{req}$ updates the world state by updating the Description Hash, User Sign, Owner Sign, BF, and Status. Since the world state is public to all participants on the consortium chain, we use the hash of the prompt Description Hash in $Tx_{req}$ to prevent the prompt from being copied or maliciously used by an attacker. The specific process of this stage is as follows:
  • User computes $\sigma_{1}=sig_{sk_{User}}(Args)$ and sends $\sigma_{1}$ along with $Args$ to Provider for content generation.
  • Provider verifies $\sigma_{1}$ with $pk_{User}$ of User. The successful verification suggests that the prompt $Args$ aligns with the intent of the User and has not been altered during transmission.
  • Provider generates $\sigma_{2}=sig_{sk_{Provider}}(Args)$ and sends it back to User. This indicates that Provider acknowledges the legitimacy of the received $Args$, enabling further content generation operations.
  • User selects a random hash function $H$ and computes Description $\textit{Description ID}=H(Args)$.
  • User verifies the legality of $\sigma_{2}$; a successful validation implies the formation of consensus between User and Provider.
  • User generates signature Owner Sign with Description Hash.
  • A transaction $Tx_{req}$ is proposed by User to update the world state with $key=\textit{Product ID}$, where Status is used to indicate the current state of P, which is prepared.
  • User retrieves the transaction ID $Txid_{req}$ from transaction $Tx_{req}$ and may choose to store it either locally or within a decentralized storage system such as IPFS [^5] https://www.ipfs.tech, which facilitates the simplification of the copyright management process.
@@
  Send $\sigma_{3}$ and generated product P back to User;
  User:
@@
  Verify the legality of $\sigma_{4}$ with $pk_{User}$;
@@
3) Data Uploading: Provider generates product P with the AIGC model based on the prompt $Args$ provided by User. Subsequently, Provider transmits P to User and initiates a transaction $Tx_{gen}$ to update additional information related to the generation process in the world state. Upon receiving product P, User initiates transaction $Tx_{upload}$ to add product P into the world state. Fig. 6 illustrates the alteration in the world state following the execution of transactions $Tx_{gen}$ and $Tx_{upload}$. In this process, the use of Product Hash in place of the original product P is aimed at preventing potential copyright infringement or misuse, while also ensuring the security of the data in AIGC-Chain.
@@
  • Provider sends both $\sigma_{3}$ and the generated product P back to User for verification.
  • User authenticates the validity of $\sigma_{3}$; only upon successful verification will User accept product P and proceed to sign it as $\sigma_{4}=sig_{sk_{User}}(P)$, which will be sent back to Provider.
  • Provider verifies the legality of $\sigma_{4}$; a successful validation implies the formation of consensus between User and Provider.
@@
  • User computes Product $\textit{Product Hash}=H(P)$ for product P.
  • A transaction $Tx_{upload}$ is proposed by User to update the world state, where Status has transitioned from generating to generated.
  • User and Provider retrieve transaction ID $Txid_{gen}$ and $Txid_{upload}$ from the blockchain respectively and can choose to store them either locally or within a decentralized storage system such as IPFS.
@@
5) Copyright Management: For a copyright dispute, Auditor can retrieve Product ID of the disputed product P by consulting the relevant world state on the AIGC-Chain. Provider and Provider are required to provide proof to Auditor to establish their connection with the disputed product. Auditor will then assess the ownership of the copyright based on the proof gathered.
  • User submits the proof $\varphi$ to Auditor, where $\varphi=(pk_{User},Txid_{req},Txid_{upload},prompt)$.
@@
  • Auditor utilizes $pk_{User}$ and $pk_{Provider}$ to verify User Sign and Model Sign retrieved from the world state of the AIGC-Chain for Product ID, to assess the correctness of the identities associated with the product-related roles.
  • Auditor employs CTrace.Check to verify the correctness of $Txid_{req}$, $Txid_{upload}$, and $Txid_{gen}$. An output of 1 indicates that the transaction is indeed related to the generation process of the AIGC product.
  • Auditor conducts an in-depth manual review of the originality, compliance, ownership, and value of the AIGC product based on the information provided by User and Provider, as well as the product information stored on AIGC-Chain.